The difference of two numbers is 91. The second is 5 less than 5 times the first. What are the two numbers?

Answer by ptaylor(2198)

Let x=first number
Then (5x-5)=second number
Now we are told that the difference between these two numbers is 91. So:
(5x-5)-x=91 get rid of parens
5x-5-x=91 add 5 to both sides
5x-5+5-x=91+5 collect like terms
4x=96 divide both sides by 4
x=24-------------------------------------first number
5x-5=5(24)-5=120-5=115---------------------------second number
ck
115-24=91
91=91
and
five less than 5 times the first=115


Hope this helps----------------ptaylor